# A Survey on Efficient Vision-Language-Action Models

### v1

@@ 3 Efficient Model Design
Advancing VLAs toward ubiquitous deployment demands balancing their superior multimodal representational capabilities with resource constraints in real-world applications, where computational efficiency remains a critical bottleneck. While VLAs have revolutionized embodied AI through seamless fusion of perception and action reasoning, their reliance on billion-parameter architectures imposes excessive inference latencies and resource demands, thwarting integration into constrained robotic systems. Efficient model design bridges this gap through lightweight yet powerful architectures and advanced compression techniques that uphold semantic integrity while sharply reducing parameter counts and computational costs. This chapter introduces a unified taxonomy comprising two complementary facets: (1) Efficient Architectures that optimizes structures via modular designs, inference accelerations, and sparse processing to enhance efficiency without compromising model potency; and (2) Model Compression strategies that minimize representational redundancy through layer pruning, quantization, and token optimization to yield compact, performant models.
@@ 3.1 Efficient Architectures
Efficient architectures underpin scalable VLAs, curtailing computational overheads through targeted innovations that sustain multimodal synergy and action acuity amid resource scarcity. As illustrated in Fig.˜5, this subsection thoroughly examines seminal contributions, from efficient attention and transformer alternatives to decoding accelerations via parallelization and generative paradigms, lightweight components, mixture-of-experts, and hierarchical processing. All representative works on efficient architectures are summarized in Tab.˜II.
@@ 3.2 Model Compression
At the forefront of resource-constrained deployment, model compression stands as the cornerstone for efficient VLAs, masterfully distilling voluminous parameter ensembles into streamlined architectures while skillfully enhancing the inference efficiency of VLAs. As illustrated in Fig.˜6, this subsection systematically surveys pivotal techniques, from layer pruning and quantization to token optimization, elucidating their general schemes and how they benefit efficiency to cultivate robust, edge-compatible architectures. Tab.˜III summarizes the representative model compression methods.
@@ 4 Efficient Training
Vision-Language-Action (VLA) models derive their profound capabilities from large-scale, pre-trained VLM backbones, inheriting robust visual understanding and commonsense reasoning. This foundation provides a significant advantage, enabling VLAs to interpret complex multimodal scenes and natural language instructions from the outset. However, this inheritance itself brings substantial burdens, rendering their training computationally intensive [51], time-consuming [1], and highly dependent on vast high-quality datasets [38]. Consequently, advancing efficient training methodologies has emerged as a paramount research objective, aiming to alleviate these resource demands without sacrificing model performance. As illustrated in Fig.˜7, this chapter systematically explores the spectrum of these techniques, focusing on two pivotal stages: (1) Efficient Pre-training, instilling foundational action capabilities into pre-trained VLMs or training the entire VLA from scratch with minimal overhead, and (2) Efficient Post-training, enabling the swift and effective deployment of VLAs to specific downstream tasks.
A VLA-centric Taxonomy of Training. Our taxonomy is defined from a functional, VLA-centric perspective. We define Pre-training as the entire process of migrating a general-purpose VLM into the embodied domain to create an initial, action-aware policy. This foundational stage is concerned with instilling the fundamental capability to act. In contrast, Post-training focuses on the subsequent specialization of this generalist policy, adapting it to excel in specific tasks, environments, or under particular resource constraints. This perspective clarifies why our pre-training VLA stage encompasses a broad spectrum of methods, including some techniques often labeled as “post-training” in other fields. The key distinction is not the method itself, but its objective: if the goal is to create the first version of the action-capable model from a VLM, we classify it as pre-training. This approach offers a more fundamental and logically consistent framework that aligns with the practical development lifecycle of VLAs.
@@ 4.1 Efficient Pre-Training
Efficient pre-training is a critical research direction aimed at decoupling VLA model performance gains from the prohibitive computational footprint incurred by large-scale foundation backbones and massive multimodal datasets. As illustrated in Fig.˜7 left, this section systematically examines advancements categorized into data-efficient pre-training, efficient action representation, and other efficient pre-training strategies. By highlighting these innovations, we outline the critical pathways for developing robust, general VLA policies without incurring the traditional burden of extensive, resource-intensive pre-training cycles. We list representative works in Tab.˜IV.
@@ 4.1.2 Efficient Action Representation
4.1.3 Other Pre-training Strategies
4.2 Efficient Post-Training
Following the foundational pre-training phase, efficient post-training techniques refine VLAs to align with downstream tasks while minimizing computational demands. As illustrated in Fig.˜7 right, these methods encompass supervised fine-tuning for targeted parameter updates and reinforcement learning-based approaches for policy optimization. Notably, efficient post-training is the crucial adaptation phase that ensures VLAs are not only specialized but also computationally viable for robust deployment in a wide array of practical applications. Tab.˜V summarize representative works.
@@ 4.2.2 RL-Based Method
5 Efficient Data Collection
5.1 Human-in-the-Loop Data Collection
5.2 Simulation Data Collection
@@ 8 Conclusion
This survey systematically charts the emergent field of Efficient VLAs, providing the first unified taxonomy focusing on the data-model-training loop to consolidate the fragmented efforts addressing their prohibitive computational and data bottlenecks. We demonstrate that current solutions converge upon three interdependent pillars, including Efficient Model Design, Efficient Training, and Efficient Data Collection. Our analysis establishes a foundational reference that illuminates the critical interplay and persistent trade-offs, such as compactness versus expressivity, defining this frontier. The challenges and future directions delineated are a direct consequence of this synthesis, charting a necessary roadmap that moves beyond isolated optimizations toward adaptive, co-designed systems. This work thus serves to catalyze this transition, accelerating the vital shift from resource-bound prototypes to truly ubiquitous physical-world intelligence.

### v2

@@ 3 Efficient Model Design
Efficient model design bridges the gap between foundational complexity and resource-constrained deployment through lightweight yet powerful architectures and advanced compression techniques that uphold semantic integrity while sharply reducing parameter counts and computational costs. This chapter introduces a unified taxonomy comprising two complementary facets of efficient model design: (1) Efficient Architectures that optimizes structures via modular designs, inference accelerations, and sparse processing to enhance efficiency without compromising model potency; and (2) Model Compression strategies that minimize representational redundancy through layer pruning, quantization, and token optimization to yield compact, performant models.
@@ 3.1 Efficient Architectures
As illustrated in Fig.˜5, this subsection thoroughly examines seminal contributions, from efficient attention and transformer alternatives to decoding accelerations via parallelization and generative paradigms, lightweight components, mixture-of-experts, and hierarchical processing. All representative works on efficient architectures are summarized in Tab.˜II.
@@ 3.2 Model Compression
As illustrated in Fig.˜6, this subsection systematically surveys pivotal techniques, from layer pruning and quantization to token optimization, elucidating their general schemes and how they benefit efficiency to cultivate robust, edge-compatible architectures. Tab.˜III summarizes the representative model compression methods.
@@ 3.2.3 Token Optimization
3.3 Discussion
3.3.1 Innovations
Efficient VLA design has transitioned from scaling-centric paradigms to adaptive architectures that integrate linear-time attention and hierarchical decoupling to alleviate inherent computational bottlenecks. This evolution, bolstered by compression strategies such as pruning and quantization, prioritizes the retention of perceptual-motor invariants through task-aware distillation. Such advancements effectively reconcile high-level reasoning with reactive execution, enabling the deployment of foundational intelligence within the stringent resource constraints of physical agents.
3.3.2 Limitations
Nevertheless, aggressive structural optimization often engenders a tension between architectural parsimony and representational fidelity, potentially inducing semantic drift that undermines dexterity in long-horizon tasks. Hierarchical and parallelized frameworks also struggle with spatiotemporal coherence, where asynchronous execution leads to unstable control signals. Furthermore, the reliance on static importance metrics limits environmental adaptability, necessitating future research into hardware-algorithm co-design to optimize the Pareto frontier of efficiency and robustness.
@@ 4 Efficient Training
While foundational VLAs leverage pre-trained VLM backbones for robust multimodal reasoning, they inherit prohibitive computational and data overheads, necessitating the advancement of efficient training methodologies to alleviate resource demands without sacrificing model performance. As illustrated in Fig.˜7, this chapter systematically explores the spectrum of these techniques, focusing on two pivotal stages: (1) Efficient Pre-training, instilling foundational action capabilities into pre-trained VLMs or training the entire VLA from scratch with minimal overhead, and (2) Efficient Post-training, enabling the swift and effective deployment of VLAs to specific downstream tasks.
A VLA-centric Taxonomy of Training. Our taxonomy is defined from a functional, VLA-centric perspective. We define Pre-training as the entire process of migrating a general-purpose VLM into the embodied domain to create an initial, action-aware policy. This foundational stage is concerned with instilling the fundamental capability to act. In contrast, Post-training focuses on the subsequent specialization of this generalist policy, adapting it to excel in specific tasks, environments, or under particular resource constraints. This perspective clarifies why our pre-training VLA stage encompasses a broad spectrum of methods, including some techniques often labeled as “post-training” in other fields [51, 93, 170, 178]. The key distinction is not the method itself, but its objective: if the goal is to create the first version of the action-capable model from a VLM, we classify it as pre-training. This approach offers a more fundamental and logically consistent framework that aligns with the practical development lifecycle of VLAs.
@@ 4.1 Efficient Pre-Training
As illustrated in Fig.˜7 left, this section systematically examines advancements categorized into data-efficient pre-training, efficient action representation, and other efficient pre-training strategies. By highlighting these innovations, we outline the critical pathways for developing robust, general VLA policies without incurring the traditional burden of extensive, resource-intensive pre-training cycles. We list representative works in Tab.˜IV.
@@ 4.1.2 Efficient Action Representation
4.1.3 Other Pre-training Strategies
4.2 Efficient Post-Training
As illustrated in Fig.˜7 right, these methods encompass supervised fine-tuning for targeted parameter updates and reinforcement learning-based approaches for policy optimization. Notably, efficient post-training is the crucial adaptation phase that ensures VLAs are not only specialized but also computationally viable for robust deployment in a wide array of practical applications. Tab.˜V summarize representative works.
@@ 4.2.2 RL-Based Method
4.3 Discussion
4.3.1 Innovations
VLA training has shifted toward modular, data-thrifty regimes that facilitate the seamless migration of multimodal priors into the embodied domain. By leveraging unlabeled egocentric videos, current methodologies instill foundational motor priors while circumventing the prohibitive costs of expert demonstrations. This process is further refined by latent action modeling and spectral compression, which distill high-dimensional control into compact manifolds to streamline the policy search space. Ultimately, decoupling foundational pre-training from task-specific refinement—particularly through reinforcement learning—enables models to achieve fine-grained dexterity without eroding the reasoning depth of their vision-language origins.
4.3.2 Limitations
Nevertheless, the kinematic discrepancy between human-centric video and robotic platforms introduces significant noise, hindering the acquisition of high-fidelity action priors. While parameter-efficient fine-tuning curtails computational overhead, it may constrain the representational flexibility required for the deep cognitive shifts essential in multi-stage manipulation. Furthermore, RL-based adaptation remains susceptible to reward misspecification and distribution shifts, where offline methods suffer from value overestimation and online paradigms face severe sample-efficiency bottlenecks. Addressing these challenges necessitates more robust alignment objectives that reconcile rapid task specialization with the preservation of multimodal coherence in unscripted environments.
5 Efficient Data Collection
5.1 Human-in-the-Loop Data Collection
5.2 Simulation Data Collection
@@ 5.5 Data Augmentation
5.6 Discussion
5.6.1 Innovations
The trajectory of VLA data acquisition has pivoted from labor-intensive teleoperation toward computation-driven regimes that effectively decouple dataset expansion from human temporal constraints. Simulation frameworks and autonomous exploration now facilitate parallelized generation, substituting manual effort with computational throughput to achieve unprecedented scale. This shift is further augmented by cross-domain manifold alignment, which leverages internet-scale egocentric videos as kinematic proxies through advanced retargeting and generative synthesis. Consequently, the integration of generative augmentation transforms the data pipeline from passive accumulation into an active process that maximizes the information density of sparse expert signals.
5.6.2 Limitations
Notwithstanding these advances, substantial bottlenecks persist in reconciling high-throughput data with real-world physical utility. The reliance on synthetic and cross-modal sources introduces inherent sim-to-real and embodiment discrepancies, where misalignments in contact dynamics and kinematic topology degrade policy transferability. While autonomous exploration alleviates data scarcity, it necessitates a delicate balance between sample efficiency and safety, often constrained by world models prone to physics hallucinations. Ultimately, the lack of precise action labels in internet-scale datasets imposes a fundamental ceiling on control fidelity, which necessitates continued reliance on in-domain expert supervision for high-precision tasks.
@@ 8 Conclusion
This survey systematically charts the emergent field of Efficient VLAs, providing the first unified taxonomy focusing on the model-training-data loop to consolidate the fragmented efforts addressing their prohibitive computational and data bottlenecks. We demonstrate that current solutions converge upon three interdependent pillars, including Efficient Model Design, Efficient Training, and Efficient Data Collection. Our synthesis uncovers the intrinsic trade-offs between model compactness and multimodal expressivity, providing a rigorous framework to navigate the systemic tensions that characterize this design space. The challenges and future directions delineated are a direct consequence of this synthesis, charting a necessary roadmap that moves beyond isolated optimizations toward adaptive, co-designed systems. This work thus serves to catalyze this transition, accelerating the vital shift from resource-bound prototypes to truly ubiquitous physical-world intelligence.